# Shortcut to Chemically Accurate Quantum Computing via Density-based Basis-set Correction

### v1

@@ Abstract
Quantum computing promises a computational advantage over classical methods in electronic-structure calculations, with expected applications in drug design and materials science. Accessing a quantitative description of chemical systems while minimizing quantum resources, such as the number of qubits, is an essential challenge given the limited capabilities of current quantum processors. We provide a shortcut towards quantum computations at chemical accuracy by approaching the complete-basis-set limit (CBS) through integrating density-functional theory into quantum algorithms via density-based basis-set corrections coupled to basis-sets crafted on-the-fly and specifically adapted to a given system/user-defined qubit budget. The approach self-consistently accelerates the basis-set convergence, improving electronic densities, ground-state energies, and first-order properties such as dipole moments. It can also serve as a classical, a posteriori, energy correction to quantum hardware calculations. The strategy is assessed using GPU-accelerated state-vector emulation up to 32 qubits. We converge the ground-state energies of four systems (He, Be, H2, LiH) within chemical accuracy of the CBS full-configuration-interaction reference, while offering a systematic increase of accuracy beyond a double-zeta quality for various molecules up to the H8 hydrogen chain. We also obtain dissociation curves for H2 and LiH that reach the CBS limit whereas for the challenging simulation of the N2 triple-bond breaking, we achieve a near-triple-zeta quality at the cost of a minimal basis-set. This hybrid strategy allows us to obtain quantitative results that would otherwise require brute-force quantum simulations using far more than 100 logical qubits, thereby opening up opportunities to explore real-world chemistry with reasonable computational resources.
@@
Quantum computing (QC) offers a promising approach to solving electronic-structure problems, with algorithms like Quantum Phase Estimation (QPE) and Variational Quantum Eigensolver (VQE) proving effective in performing ground-state quantum-chemistry wave-function calculations [1, 2, 3, 4, 5, 6]. The electronic Hamiltonian is expressed in second quantization, employing an encoding that maps one spin-orbital to one qubit. This mapping allows one to represent an exponentially large Hilbert space using only a linear number of qubits. However, to achieve accurate and practically valuable predictions for chemical systems in real-world applications, the molecular Hamiltonian should be expressed and solved using extensive basis sets $\mathcal{B}$ of orbital functions. The number of qubits required for such calculations quickly exceeds the available capacities on current Noisy Intermediate Scale Quantum (NISQ) devices, upcoming early Fault-Tolerant Quantum Computing (FTQC) devices, and high-performance classical emulators. Therefore, while chemistry has long been stated as a promising application for quantum computing, the endeavors have so far been limited to small molecular systems and minimal basis sets. However, in practice, minimal basis sets fail to be predictive for ground-state energies and chemically useful calculations require at least significantly larger double-zeta basis sets. Requirements for computing molecular properties are even more drastic as they tend to converge more slowly with the size of the basis set than energies [7].
In electronic-structure theory, the exact solution is defined by the full-configuration interaction (FCI) method in the (infinite) complete-basis-set (CBS) limit. However, in practice, the solution derived from a finite basis set $\mathcal{B}$ is used, which inherently suffers from truncation errors. This errors can be substantial for small basis sets, but for large enough basis sets it is possible to reach chemical accuracy, 1 kcal/mol (1.6 mHa), which is actually the target. Unfortunately, employing sufficiently large basis sets becomes very rapidly impractical for large systems, particularly those with more than a few dozen of light atoms. As a result, traditional quantum chemistry has developed a variety of methods to approach chemical accuracy at a reasonable cost.
In particular, basis-set correction techniques such as the density-based basis-set correction (DBBSC) method [8, 7, 9, 10, 11, 12, 13, 14, 15, 16], relying on density-functional theory (DFT), have proven effective not only for calculating ground- and excited-state energies, but also for determining dipole moments for a variety of systems including atoms, small organic molecules, and strongly correlated systems. To date, the DBBSC method has only been applied to still relatively large basis sets, and more specifically the family of Dunning basis sets [17], where chemical accuracy with respect to the CBS limit can be reached starting from a triple-zeta basis set. We note that the approximations developed within the DBBSC method target the CBS limit within a wave-function approach. This means that they do not address the intrinsic errors of the wave-function method itself, such as the effects of neglected excitation terms in truncated configuration-interaction (CI) or coupled-cluster methods.
@@
Figure 1: Architecture of the protocol introducing the DBBSC method to quantum computing. Within this hybrid quantum-classical computing scheme, the QPU (Quantum Processing Unit) workload can be either performed by quantum hardware or replaced by a GPU-accelerated quantum emulator (GPU= Graphics Processing Unit, see section I.C.). The system and basis-set information are encoded in the quantum state. In Strategy 1, the QPU/GPU computes $\langle H\rangle$ and the non-self-consistent correlation correction and the HF basis-set correction are shifted to the CPUs (Central Processing Unit) of the classical compute node. In Strategy 2, the model includes a self-consistent basis-set correction potential which is iteratively optimized between QPU/GPU and CPU. The insets depict the effect of the corrections for the two strategies. One can choose either standard basis sets or SABS.
@@
Here, we propose the integration of the DBBSC method with quantum algorithms to expedite reaching the CBS limit and achieve chemical accuracy on large complex molecular systems. This strategy, which natively limits the required qubit counts, can be applied to quantum algorithms that tackle the ground-state quantum chemistry problem, such as the QPE or the VQE algorithms. While QPE can guarantee ground-state energy with arbitrary high precision given a carefully chosen initial state, it demands large quantum circuits and will only be viable in the FTQC era. Conversely, VQE lacks convergence guarantees but employs smaller circuits, aligning better with this study’s aim of advancing short-term quantum computers toward practical quantum chemistry applications.
@@
In section I, we present two approaches (denoted ”Strategy 1” and ”Strategy 2”): 1) a basis-set correction a posterioriadded to the solution of the quantum algorithm, integrating two contributions, namely the DBBSC correction developed in previous work [8] and a correction to the Hartree-Fock (HF) energy, and: 2) a self-consistent scheme [12] integrating the DBBSC method to the quantum algorithm that dynamically modifies the density-dependent terms in the Hamiltonian. Strategy 1 offers the possibility to correct any QC energy computation through an additive correction. Strategy 2 enables to self-consistently access an improved electronic density, offering improved energies and access to first-order properties.
@@
These new methodologies enable the first investigation of DBBSC corrections in the minimal basis-set regime. To enhance further the accuracy of the DBBSC method beyond minimal basis sets, it is always possible to choose a larger basis set. In section II, we provide relevant tests using the main families of basis sets in Table 2. Alternatively, in order to limit the qubit counts and to avoid the basis-set choice (which can be sometimes troublesome), we introduce a new strategy for defining compact adaptive basis sets . This on-the-fly approach is grounded on a pivoted Cholesky approach. We carried out numerical simulations using GPU-accelerated QC sparse emulation on up to 32 qubits, exploring the applicability of this method to ground-state energies, dissociation curves, and dipole moments. We consistently observed significant improvements over typical quantum-algorithm approaches, reaching an accuracy level that would have otherwise required hundreds of qubits. We expect this to become part of a suite of ansatz improvements that will be employed in quantum-enhanced chemistry calculations while the two strategies based on the DBBSC method, particularly Strategy 1, can immediately provide improvements to existing results, with relatively simple effort.
@@ I Methodology
The overall procedure is illustrated in Figure 1. The procedure starts with the definition of the system and basis set such as STO-3G, 6-31G, or pcseg-0 [18]. Also, we introduce a new type of system-adapted basis sets (SABS) for Gaussian-type orbitals (GTOs), with size comparable to minimal basis sets, generated by taking as input the HF one-electron density of the system of interest at a fixed geometry (geometries and SABS used in this paper are available in the Supplementary Information (SI), see section VI A. and VI C.).
To build a SABS, the one-electron density is expanded on an accurate basis set, such as a Dunning basis set where the required number of qubits can reach the hundreds. This generation is based on the pivoted Cholesky decomposition (see mathematical details in Section I.4), which is a theoretical tool commonly used to eliminate linear dependencies from a given basis set in ab initio electronic-structure theory [19, 20, 21]. In Table 2 and 3, the SABS are labeled VXZ-Y where X is respectively D, T, Q, 5, or 6 for cc-pVDZ, cc-pVTZ, up to cc-pV6Z basis sets, and Y is the target size. In practice, the generated basis can have a larger size due to some additional constraints such as the need to keep the same basis functions for a given atom type particularly when spatial symmetries occur. We report in the SI the classically computed energies for the N2, H2O, LiH, and H2 molecules.
We note that the adaptation and generation of basis sets to the molecular geometry has been the subject of several publications exploring other strategies. Among them, we can cite some recent works [22, 23], but also some other research in the context of quantum computing focusing on the need to limit the qubit requirements through basis-set reoptimization such as Refs. [24, 25]. Such reoptimizations cannot be performed on the fly and obtaining such improvements can take a large amount of time when dealing with large basis sets. Alternatively, Kottman and Aspuru [26] proposed a basis-set-free approach through an adaptive representation using pair-natural orbitals which was tested up to 22 qubits and potentially able to deal with systems in the range of 40 to 100 qubits.
After defining the second-quantized Hamiltonian, the quantum state is prepared on a quantum processing unit (QPU), using standard Hamiltonian mapping as Jordan-Wigner and Bravyi-Kitaev transformations (see SI, section II.). This step is followed by one of the two procedures labeled as “Strategy 1” and “Strategy 2”.
I.1 Strategy 1: a posteriori basis-set correction
“Strategy 1” relies on independent contributions of the quantum and classical computation parts, allowing applications to non-variational quantum algorithms. It can be used as an affordable and direct correction to any QC computation performed on real quantum hardware. The non-self-consistent correlation basis-set correction $\Delta E_{1}^{\mathcal{B},\text{c}}$ and the HF basis-set correction $\Delta E_{1}^{\mathcal{B},\text{HF}}$ are computed on classical computing units and added to the expectation value $E_{1}^{\mathcal{B}}$ coming out of the quantum solver. Therefore, the basis-set corrected energy in the basis set $\mathcal{B}$ is approximated as
$E_{0}^{\mathcal{B}}\simeq E^{\mathcal{B}}_{1}+\Delta E_{1}^{\mathcal{B},\text{c}}+\Delta E_{1}^{\mathcal{B},\text{HF}}.$ (1)
We explicit the correction quantities in the following paragraphs and further details can be found in the SI (see Section III).
I.1.1 Non-self-consistent correlation basis-set correction
The non-self-consistent correlation basis-set correction is a functional of the one-electron density $n^{\mathcal{B}}$,
$\Delta E_{1}^{\mathcal{B},\text{c}}=\bar{E}^{\mathcal{B}}[n^{\mathcal{B}}],$ (2)
and is built to correct the basis-set incompleteness due to short-range electron correlation in such a way that the correction correctly vanishes in the CBS limit: $\lim_{\mathcal{B}\to\text{CBS}}\bar{E}^{\mathcal{B}}[n^{\mathcal{B}}]=0$. In this work, we use the approximation based on the Perdew-Burke-Ernzerhof (PBE) correlation functional introduced in Ref. [7]. We also discuss the choice of the approximation in the SI which we advise the reader to keep for a second reading as it does not add to the understanding of the procedure. However, we can mention here that the use of a better functional could give access to stronger-correlation regimes and further work on alternative approximations of the functional could allow applications on more complex systems. As in the initial work of the DBBSC method [8], we use the HF density $n_{\text{HF}}^{\mathcal{B}}$ as an input density, leading to a correction with marginal computational cost with respect to the cost of calculating $E_{1}^{\mathcal{B}}$.
I.1.2 HF basis-set correction
The HF basis-set correction $\Delta E_{1}^{\mathcal{B},\text{HF}}$ corrects for the error in the HF energy occurring using a small basis set. Initial developments of the DBBSC approach were applied to relatively large basis sets where the HF energy has a maximum error of 0.1 mHa. For its application to quantum computations where minimal basis sets are usually chosen, the error within the HF energy is comparable to the electronic correlation error. However, this error is not taken into account in the DBBSC correlation approximations. A direct way to evaluate the HF basis-set correction $\Delta{E}_{1}^{\mathcal{B},\text{HF}}$ is to compute the difference between the HF energy in the CBS limit, that we estimate as the value $E_{\text{HF}}^{\text{5Z}}$ obtained with cc-pV5Z basis set, and the HF energy $E_{\text{HF}}^{\mathcal{B}}$ in the basis set $\mathcal{B}$
$\Delta E_{1}^{\mathcal{B},\text{HF}}=E_{\text{HF}}^{\text{5Z}}-E_{\text{HF}}^{\mathcal{B}}.$ (3)
I.2 Strategy 2: Self-consistent basis-set correction
In “Strategy 2”, we use the self-consistent version of the DBBSC method[8, 12]. In the approach, the Hamiltonian used in the quantum solver is
$\hat{\tilde{H}}=\hat{H}+\hat{\bar{V}}[n^{\mathcal{B}}],$ (4)
where
$\hat{H}=\hat{T}+\hat{V}_{\text{ne}}+\hat{W}_{\text{ee}},$ (5)
where $\hat{T}$, $\hat{V}_{\text{ne}}$, and $\hat{W}_{\text{ee}}$ are respectively the kinetic, nuclei-electron interaction, and electron-electron interaction operators, and $\hat{\bar{V}}^{\mathcal{B}}[n^{\mathcal{B}}]$ is the one-electron basis-set correction potential operator,
$\hat{\bar{V}}[n^{\mathcal{B}}]=\int\bar{v}^{\mathcal{B}}[n^{\mathcal{B}}](\textbf{r})\hat{n}^{\mathcal{B}}(\textbf{r})\text{d}\textbf{r},$ (6)

### v2

@@ Abstract
Quantum computing promises a computational advantage over classical methods in electronic-structure calculations, with expected applications in drug design and materials science. Accessing a quantitative description of chemical systems while minimizing quantum resources, such as the number of qubits, is an essential challenge given the limited capabilities of current quantum processors. We provide a shortcut towards quantum computations at chemical accuracy by approaching the complete-basis-set (CBS) limit through integrating density-functional theory into quantum algorithms via density-based basis-set corrections coupled to basis-sets crafted on-the-fly and specifically adapted to a given system/user-defined qubit budget. The approach self-consistently accelerates the basis-set convergence, improving electronic densities, ground-state energies, and first-order properties such as dipole moments. It can also serve as a classical, a posteriori, energy correction to quantum hardware calculations. The strategy is assessed using GPU-accelerated state-vector emulation up to 32 qubits. We converge the ground-state energies of four atomic and molecular systems (He, Be, H2, LiH) within chemical accuracy of the CBS full-configuration-interaction reference, while offering a systematic increase of accuracy beyond a double-zeta basis-set quality for various molecules up to the H8 hydrogen chain. We also obtain dissociation curves for H2 and LiH that reach the CBS limit whereas for the challenging simulation of the N2 triple-bond breaking we achieve a near-triple-zeta basis-set quality at the cost of a minimal basis set. This hybrid strategy allows us to obtain quantitative results that would otherwise require brute-force quantum calculations using far more than 100 logical qubits, thereby opening up opportunities to explore real-world chemistry with reasonable computational resources.
@@
Quantum computing (QC) offers a promising approach to solving electronic-structure problems, with algorithms like Quantum Phase Estimation (QPE) and Variational Quantum Eigensolver (VQE) proving effective in performing ground-state quantum-chemistry wave-function calculations [1, 2, 3, 4, 5, 6]. The electronic Hamiltonian is expressed in second quantization, employing an encoding that maps one spin-orbital to one qubit. This mapping allows one to represent an exponentially large Hilbert space using only a linear number of qubits. However, to achieve accurate and practically valuable predictions for chemical systems in real-world applications, the molecular Hamiltonian should be expressed and solved using extensive basis sets of one-electron orbital functions. The number of qubits required for such calculations quickly exceeds the available capacities on current Noisy Intermediate Scale Quantum (NISQ) devices, upcoming early Fault-Tolerant Quantum Computing (FTQC) devices, and high-performance classical emulators. Therefore, while quantum chemistry has long been stated as a promising application for quantum computing, the endeavors have so far been limited to small molecular systems and minimal basis sets. However, in practice, minimal basis sets fail to be predictive for ground-state energies and chemically useful calculations require at least significantly larger double-zeta basis sets. Requirements for computing molecular properties are even more drastic as they tend to converge more slowly with the size of the basis set than energies.
@@
Figure 1: Architecture of the hybrid quantum-classical scheme introducing the DBBSC method to wave-function QC calculations. The QPU workload can be either performed by quantum hardware or replaced by a GPU-accelerated quantum emulator. The input are the definition of the system and the basis set (standard basis sets or our SABS can be used). In Strategy 1, the QPU/GPU computes the expectation value of the standard Hamiltonian $\langle\hat{H}\rangle$ and the non-self-consistent correlation basis-set correction and the HF basis-set correction are shifted to the CPU of the classical computer. In Strategy 2, the model includes a self-consistent basis-set correction potential which is iteratively optimized between QPU/GPU and CPU. The insets depict the effect of the basis-set corrections for the two strategies.
In electronic-structure theory, the exact solution is defined by the full-configuration interaction (FCI) method in the (infinite) complete-basis-set (CBS) limit. However, in practice, the solution derived from a finite basis set is used, which inherently suffers from truncation errors. This errors can be substantial for small basis sets, but for large enough basis sets it is possible to reach the target of chemical accuracy on energy differences, i.e. 1 kcal/mol (1.6 mHa). Unfortunately, employing sufficiently large basis sets becomes very rapidly impractical for large systems, particularly those with more than a few dozen of atoms. As a result, traditional quantum chemistry has developed a variety of computational methods to approach chemical accuracy at a reasonable cost.
In particular, basis-set correction techniques such as the density-based basis-set correction (DBBSC) method [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], relying on density-functional theory (DFT), have proven effective for calculating ground- and excited-state energies, and also dipole moments, for a variety of systems including atoms, small organic molecules, and strongly correlated systems. To date, the DBBSC method has only been applied to still relatively large basis sets, more specifically the family of Dunning basis sets [20] for which chemical accuracy with respect to the CBS limit can be reached starting from a triple-zeta basis set. We note that the approximations developed within the DBBSC method target the CBS limit within a given wave-function ansatz. This means that they do not address the intrinsic errors of the wave-function method itself, such as the effects of neglected higher excitations in truncated configuration-interaction (CI) or coupled-cluster methods.
@@
In this work, we propose the integration of the DBBSC method with quantum algorithms to expedite reaching the CBS limit and achieve chemical accuracy on complex molecular systems. This strategy, which natively limits the required qubit counts, can be applied to quantum algorithms that tackle the ground-state quantum chemistry problem, such as the QPE or VQE algorithms. While QPE can guarantee ground-state energy with arbitrary high precision given a carefully chosen initial state, it demands large quantum circuits and will only be viable in the FTQC era. Conversely, VQE lacks convergence guarantees but employs smaller circuits, aligning better with this study’s aim of advancing short-term quantum computers toward practical quantum-chemistry applications.
@@
The paper is organized as follows. In Section I, we present two variants of the application of the DBBSC method to wave-function QC calculations, denoted as Strategy 1 and Strategy 2: (1) a basis-set correction a posteriori added to the solution of the quantum algorithm, integrating two contributions, namely a basis-set correlation density-functional correction and a basis-set Hartree-Fock (HF) correction, and (2) a self-consistent scheme integrating the DBBSC method to the quantum algorithm that dynamically modifies the one-electron density used in the basis-set correlation correction. Strategy 1 offers the possibility to correct any wave-function QC energy calculation through a simple additive correction. Strategy 2 enables one to self-consistently access to an improved electronic density, offering both improved energies and first-order molecular properties. Also, we introduce a new type of system-adapted basis sets (SABS) for Gaussian-type orbitals (GTOs), with size comparable to minimal basis sets. These new methodologies enable us to perform the first investigation of DBBSC corrections in the minimal basis-set regime.
@@
In Section II, we provide relevant tests of our approach for different atomic and molecular systems using several families of basis sets. We carry out numerical simulations using graphics-processing-unit (GPU) accelerated QC sparse emulation on up to 32 qubits, exploring the applicability of this method to converge ground-state energies, dissociation curves, and dipole moments. We consistently observe significant improvements over typical quantum-algorithm approaches, reaching an accuracy level that would have otherwise required hundreds of qubits. We thus expect the present approach to become a standard part of quantum-enhanced wave-function calculations, particularly the approach of Strategy 1 which can immediately provide large improvements to existing results with relatively simple efforts.
Finally, Section III contains our conclusions. Additional details and results are provided in the Supplementary Information (SI).
@@ I Methodology
The overall methodological procedure is illustrated in Figure 1. The procedure starts with the definition of the system, and a standard basis set or our SABS. After defining the second-quantized Hamiltonian, the quantum state is prepared on a quantum processing unit (QPU) or a GPU-accelerated quantum emulator. The basis-set corrections are calculated on the classical CPU using one of the two DBBSC variants labeled as Strategy 1 and Strategy 2.
I.1 Density-based basis-set correction method
In the infinite-dimensional (antisymmetric) $N$-electron Hilbert space, ${\cal H}=\bigwedge^{N}L^{2}(\mathbb{R}^{3},\mathbb{C})$, we consider an atomic or molecular system with Hamiltonian
$\hat{H}=\hat{T}+\hat{W}_{\text{ee}}+\hat{V}_{\text{ne}},$ (1)
where $\hat{T}$ is the kinetic-energy operator, $\hat{W}_{\text{ee}}$ is the Coulomb electron-electron operator, and $\hat{V}_{\text{ne}}$ is the nuclei-electron potential operator. The exact ground-state energy is defined as
$E_{0}=\min_{\Psi\in{\cal W}}\langle\Psi,\hat{H}\Psi\rangle,$ (2)
[…]
$E_{\text{FCI}}^{\cal B}=\min_{\Psi\in{\cal W}^{\cal B}}\langle\Psi,\hat{H}\Psi\rangle,$ (3)
where ${\cal W}^{\cal B}=\{\Psi\in{\cal H}^{\cal B}\;|\;\langle\Psi,\Psi\rangle=1\}$ is the space of ${\cal B}$-representable wave functions. In the CBS limit, the FCI ground-state energy tends to the exact ground-state energy, i.e. $E_{\text{FCI}}^{{\cal B}\to\text{CBS}}=E_{0}$, but the convergence is infamously slow due to short-range electron correlation [21].
In the DBBSC method [7, 13], for the given basis set ${\cal B}$, we introduce the following approximation to the ground-state energy
$E_{0}^{\cal B}=\min_{\Psi\in{\cal W}^{\cal B}}\left(\langle\Psi,\hat{H}\Psi\rangle+\bar{E}^{\cal B}[n_{\Psi}]\right),$ (4)
where $\bar{E}^{\cal B}[n_{\Psi}]$ is a basis-set correction density functional evaluated at the one-electron density of $\Psi$. This basis-set correction density functional must vanish in the CBS limit so that $E_{0}^{\cal B}$ properly converges to the exact ground-state energy, i.e. $E_{0}^{{\cal B}\to\text{CBS}}=E_{0}$, but it must be such that it accelerates the convergence to the CBS limit. In Ref. [8], based on the Perdew-Burke-Ernzerhof (PBE) correlation density functional [22], such a basis-set correction density functional was constructed in a semilocal form
$\bar{E}^{\cal B}[n]=\int_{\mathbb{R}^{3}}\bar{e}^{\cal B}(n(\textbf{r}),\nabla n(\textbf{r}))\text{d}\textbf{r},$ (5)
where $n(\textbf{r})$ and $\nabla n(\textbf{r})$ are the density and density gradient at point r, and the function $\bar{e}^{\cal B}(n,\nabla n)$, including how the dependence on the basis set ${\cal B}$ is included, can be found in Ref. [8]. We will now discuss how the DBBSC method can be adapted to wave-function QC calculations.
I.1.1 Strategy 1: a posteriori basis-set correction
In Strategy 1, we use a non-self-consistent approximation to Eq. (4). The idea is to calculate the FCI energy $E_{\text{FCI}}^{\cal B}$, or a good approximation to it, on a quantum computer, and then add a posteriori the basis-set correlation correction of Eq. (5) and a HF basis-set correction, both calculated on a classical computer. This leads to the following approximation to the ground-state energy
$E_{1}^{\cal B}=E_{\text{FCI}}^{\cal B}+\bar{E}^{\cal B}[n_{\text{HF}}^{\cal B}]+\Delta E_{\text{HF}}^{\cal B}.$ (6)
As in previous works [7, 8], the basis-set correlation correction is evaluated at the HF density in the basis set ${\cal B}$, leading to a calculation of $\bar{E}^{\cal B}[n_{\text{HF}}^{\cal B}]$ with a marginal computational cost with respect to the cost of calculating $E_{\text{FCI}}^{\cal B}$. The PBE-based basis-set correlation correction only corrects for basis-set incompleteness errors due to short-range electron correlation. However, when using small basis sets, a significant part of the basis-set error also comes from the fact the HF part of the energy is not converged to the CBS limit. Similarly to Refs. [18, 19], we thus add a HF basis-set correction. In the present work, we choose it simply as the difference between the HF energy in the CBS limit, that we estimate as the value $E_{\text{HF}}^{\text{5Z}}$ obtained with the cc-pV5Z basis set, and the HF energy $E_{\text{HF}}^{\mathcal{B}}$ in the basis set $\mathcal{B}$
$\Delta E^{\mathcal{B}}_{\text{HF}}=E_{\text{HF}}^{\text{5Z}}-E_{\text{HF}}^{\mathcal{B}}.$ (7)
We note that it is also possible to avoid performing a HF calculation with the cc-pV5Z basis set for estimating the CBS limit of the HF energy by using a complementary auxiliary basis set, as in Refs. [18, 19], but we have not found it necessary for the present work.
I.1.2 Strategy 2: Self-consistent basis-set correction
In Strategy 2, we use the self-consistent version of the DBBSC method in Eq. (4), using the basis-set correlation correction of Eq. (5) in the self-consistent part of the calculation, and we only add a posteriori the fixed HF basis-set correction. This leads to the following approximation to the ground-state energy
$E_{2}^{\cal B}=\min_{\Psi\in{\cal W}^{\cal B}}\left(\langle\Psi,\hat{H}\Psi\rangle+\bar{E}^{\cal B}[n_{\Psi}]\right)+\Delta E^{\mathcal{B}}_{\text{HF}}.$ (8)
The minimization in Eq. (8) leads to the following self-consistent Schrödinger equation [13]
$\hat{P}^{\cal B}\hat{\tilde{H}}^{\cal B}[n_{\Psi^{\cal B}}]\Psi^{\cal B}={\cal E}^{\cal B}\Psi^{\cal B},$ (9)
where $\hat{P}^{\cal B}$ is the projector on the $N$-electron Hilbert space generated by the basis set ${\cal B}$, i.e. ${\cal H}^{\cal B}$, and $\hat{\tilde{H}}^{\cal B}[n]$ is the effective Hamiltonian
$\hat{\tilde{H}}^{\cal B}[n]=\hat{H}+\hat{\bar{V}}^{\cal B}[n].$ (10)
Here, $\hat{\bar{V}}^{\cal B}[n]$ is the basis-set correction one-electron potential operator
$\hat{\bar{V}}^{\cal B}[n]=\int_{\mathbb{R}^{3}}\bar{v}^{\mathcal{B}}[n](\textbf{r})\;\hat{n}(\textbf{r})\text{d}\textbf{r},$ (11)